# Wenhao Wang

### v2

@@ 5.2 The Challenge from Novel Patterns
Table 2: The comparison between different visual prompting methods for incorporating image-replica pairs. Our ImageStacker not only achieves the highest performance but also maintains efficient training and inference.
[…]
@@
5.3 The Benefits from AnyPattern and In-context Learning
To improve performance on novel patterns, we show that both directly training models on larger pattern sets and conducting in-context learning are beneficial. As demonstrated in Table 1, we initially train models on the base patterns of our AnyPattern. This expansion significantly enhances performance on novel patterns: for example, resulting in a gain of $26.66\%$ in $\mu AP$ and $27.32\%$ in $recall@1$ for our baseline. However, a noticeable performance gap persists compared to scenarios where training and testing occur on the same pattern sets. This emphasizes the necessity of introducing in-context learning methods to further enhance performance. Our in-context learning method (ImageStacker) not only further improves performance on novel patterns significantly, achieving gains of $+16.75\%$ in $\mu AP$ and $+15.30\%$ in $recall@1$ for our baseline, but also maintains (marginally improves) the performance on base patterns (see Section F in the Appendix).
5.4 Comparing ImageStacker against Common Visual Prompting Methods
@@
Models can integrate image-replica pairs at both feature and image levels. At the feature level, we append an extra self-attention layer to the last block of ViT to enable interaction between the class token of the replica and the patch tokens or class token of the image-replica pairs (see Fig. 5 (2) and (3)). At the image level, besides our proposed ImageStacker, we can directly add the image-replica pairs to the replica (See Fig. 5 (4)) or concatenate the images along the height (width) dimension (see Fig. 5 (5)). We compare these different designs in Table 2, drawing three main observations: (1) Incorporating image-replica pairs at the feature level is ineffective, resulting in about a $4\sim 6\%$ performance drop in $\mu AP$. This is attributable to insufficient interaction between the replica and its image-replica pair. (2) Directly adding image-replica pairs to the replica significantly degrades the performance. This is because imposing a strong priori restriction without specific meaning can be detrimental. (3) Both ImageCombiner and ImageStacker significantly improve performance. Compared to ImageCombiner, ImageStacker achieves a greater performance gain ($+8.18\%$ in $\mu AP$ and $+7.52\%$ in $recall@1$), while requiring only about $1/4$ inference workload and $1/3$ training workload. Also, ImageStacker enhances the baseline by $13.81\%$ in $\mu AP$ and $13.00\%$ in $recall@1$, with nearly the same efficiency.
5.5 AnyPattern Enables In-context ICD
In this section, from a data perspective, we explore factors beyond our proposed ImageStacker that contribute to the emergence of in-context learning. We first adjust the training patterns of ImageStacker from AnyPattern to SmallPattern (the one used in BoT [12]), and subsequently discard the ImageNet-pre-trained models. The performance of the baseline and ImageStacker is presented in Table 3. The Upper Bound is achieved by using the query and its original image as the image-replica (example) during the inference. Our analysis reveals that: data plays a crucial role alongside the model: (1) In-context learning does not emerge when using SmallPattern and ImageNet-pretrained models: comparing against baseline with $16.18\%$ in $\mu AP$, ImageStacker only achieves $15.28\%$ in $\mu AP$. (2) The use of AnyPattern leads to the emergence of in-context learning: even without ImageNet, training ImageStacker on AnyPattern already significantly improves performance ($+13.01\%$ in $\mu AP$ and $+12.35\%$ in $recall@1$). This reaffirms the value of our proposed AnyPattern.
Beyond the above experiments, in the Appendix, we show that (1) the proposed pattern retrieval method is effective (Section G), (2) utilizing multiple image-replica pairs per query further improves the performance (Section I), and (3) our method remains effective even with fewer image-replica pairs in the pool (Section J).
Table 3: The performance of the baseline and our ImageStacker with different (pre-)training data.
@@ 5.5 AnyPattern Enables In-context ICD
$\checkmark$ $\checkmark$ $42.84$ $47.86$
Baseline $\checkmark$ $\checkmark$ $16.18$ $20.54$
$\checkmark$ $40.52$ $46.16$
$\checkmark$ $\checkmark$ $56.65$ $60.86$
ImageStacker $\checkmark$ $\checkmark$ $15.28$ $20.10$
$\checkmark$ $53.53$ $58.51$
$\checkmark$ $\checkmark$ $99.25$ $99.84$
Upper Bound $\checkmark$ $\checkmark$ $18.11$ $22.83$
$\checkmark$ $99.26$ $99.86$
@@ 6 Conclusion
This paper considers a practical scenario, i.e. in-context Image Copy Detection (ICD). Unlike the standard updating of ICD, in-context ICD aims to prompt a trained model to recognize novel-patterned replication using a few example pairs, without requiring re-training. To advance research on in-context ICD, we present AnyPattern, a dataset featuring $100$ tampering patterns. We further propose ImageStacker, a method that directly stacks an example pair onto a query along the channel dimension. The stacking design conditions (modifies) the query feature and thus enables a better matching for the novel patterns. Experimental results highlight the substantial performance improvement gained by AnyPattern and ImageStacker. We hope our work draws research attention to the critical real-world problem in ICD systems, i.e., the fast reaction against novel patterns. Beyond the problem of ICD, we also explore the value of AnyPattern in identifying style mimicry by text-to-image models. Limitations and future work. Although training ImageStacker with AnyPattern significantly improves performance on novel patterns, it still falls short compared to base patterns. Based on AnyPattern, future work may focus on developing more effective and efficient in-context learning methods to reduce overfitting on base patterns and close this performance gap.
@@ Appendix A Demonstration of the AnyPattern dataset
This section shows the $100$ patterns utilized in the creation of the AnyPattern dataset. The original image is presented in Fig. 6; the generated replicas are in the following tables, with the names of the training patterns indicated in black, and the names of the test patterns depicted in blue. The majority of the constructed patterns includes a degree of randomness while a minority of the patterns remain constant. To demonstrate the variability, each pattern is replicated four times for a single image. To our knowledge, the assembled AnyPattern is the most extensive pattern set currently available.
Figure 6: The original image that used to add $100$ different patterns
@@
Randomly shift, swap, or invert the channel of an image [11]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly enhance the edge contrast of an image, making it appear more defined and detailed. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly rearrange (shuffle) the pixels within an image, creating a significant disruption to the original structure. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Replicate the effect of a kaleidoscope on an image by creating symmetrical, repeating patterns randomly. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Simulate the effect of viewing an image through a medium that randomly distorts the light passing through it. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly simulate the effect of zooming in or out while capturing an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply a random Fourier Domain Adaptation [28] to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Dividing two random images into patches and then interweave these patches to create a new composite image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Random alter the intensities of the RGB channels in an image, based on Principal Component Analysis (PCA) of the set of RGB pixel values [29]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply random non-linear distortions within a grid to an image, producing a variety of warping effects. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
SuperPixel are the random result of perceptual grouping (image over-segmentation) of pixels in an image [30]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly change an image with RGB space to HSV, HLS, LAB, YCrCb, LUV, XYZ, or YUV space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly simulate the effect of "melting" pixels down the image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly replace the style of an image with another one’s [31]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly transfer an image to animation style [32]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly perform adversarial attacks to an image to generate adversarial example [33]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply Contrast Limited Adaptive Histogram Equalization (CLAHE) [34] to equalize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply DropOut [35] or DropBlock [36] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply WaveBlock [37] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply Wavelet [38] to an image ans select random frequency as the final result. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly change the relationship between bright and dark areas of the image by manipulating its tone curve. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Erase a random area of an image [39]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply AutoAugment [40], which is an automated data augmentation technique discovered by the reinforcement learning, to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Use the Segment Anything [41] model to preform segmentation on an image automatically. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Remove the haze from an image using dark channel prior [42]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v3

@@ 6.2 The Challenge from Novel Patterns
6.3 The Benefits from AnyPattern and In-context Learning
To improve performance on novel patterns, we show that both directly training models on larger pattern sets and conducting in-context learning are beneficial. As demonstrated in Table 1, we initially train models on the base patterns of our AnyPattern. This expansion significantly enhances performance on novel patterns: for example, resulting in a gain of $26.66\%$ in $\mu AP$ and $27.32\%$ in $recall@1$ for our baseline. However, a noticeable performance gap persists compared to scenarios where training and testing occur on the same pattern sets. This emphasizes the necessity of introducing in-context learning methods to further enhance performance. Our in-context learning method (ImageStacker) further improves performance on novel patterns significantly, achieving gains of $+16.75\%$ in $\mu AP$ and $+15.30\%$ in $recall@1$ for our baseline. Furthermore, it is also crucial to maintain performance on base patterns while enhancing performance on novel patterns through in-context learning because the edited copies generated by the base patterns may still appear in the future. As illustrated in Table 2, our proposed ImageStacker achieves this objective. For instance, with our baseline, the $\mu AP$ has been increased from $79.37\%$ to $83.56\%$, and the $recall@1$ has improved from $83.05\%$ to $85.85\%$.
Table 3: The performance of the baseline and our ImageStacker with different (pre-)training data. Without large-scale AnyPattern, in-context learning does not emerge.
  Method
AnyPattern
SmallPattern
ImageNet
$\mu AP$ $R@1$
$\checkmark$ $\checkmark$ $42.84$ $47.86$
Baseline $\checkmark$ $\checkmark$ $16.18$ $20.54$
$\checkmark$ $40.52$ $46.16$
$\checkmark$ $\checkmark$ $\mathbf{56.65}$ $\mathbf{60.86}$
Image- $\checkmark$ $\checkmark$ $15.28$ $20.10$
Stacker $\checkmark$ $53.53$ $58.51$
$\checkmark$ $\checkmark$ $99.25$ $99.84$
Upper $\checkmark$ $\checkmark$ $18.11$ $22.83$
Bound $\checkmark$ $99.26$ $99.86$
6.4 AnyPattern Enables In-context ICD
In this section, from a data perspective, we explore factors beyond our proposed ImageStacker that contribute to the emergence of in-context learning. We first adjust the training patterns of ImageStacker from AnyPattern to SmallPattern (the one used in BoT (Wang et al., 2021)), and subsequently discard the ImageNet-pre-trained models. The performance of the baseline and ImageStacker is presented in Table 3. The Upper Bound is achieved by using the query and its original image as the image-replica (example) during the inference. Our analysis reveals that: data plays a crucial role alongside the model: (1) In-context learning does not emerge when using SmallPattern and ImageNet-pretrained models: comparing against baseline with $16.18\%$ in $\mu AP$, ImageStacker only achieves $15.28\%$ in $\mu AP$. (2) The use of AnyPattern leads to the emergence of in-context learning: even without ImageNet, training ImageStacker on AnyPattern already significantly improves performance ($+13.01\%$ in $\mu AP$ and $+12.35\%$ in $recall@1$). This reaffirms the value of our proposed AnyPattern dataset.
6.5 In-context Learning Surpasses Fine-tuning
Beyond the efficiency advantage, in-context learning also offers two performance advantages, as show in Table 4: (1) Fine-tuning fails when the amount of training data is limited. For instance, when using the $1,200$ image-replica pairs ($2,400$ images) in AnyPattern for fine-tuning instead of in-context learning, the $\mu AP$ on novel patterns is $41.15\%$, which is $-15.50\%$ compared to our in-context solution; and (2) Fine-tuning on large-scale data generated by novel patterns can lead to catastrophic forgetting of the base patterns. For instance, when using $200,000$ images generated by novel patterns for fine-tuning, compared to the in-context solution, although there is a performance ($+10.87\%$ $\mu AP$) superiority on novel patterns, the performance on base patterns drops significantly ($-63.5\%$ $\mu AP$).
Table 4: The performance comparison between the in-context solution against the fine-tuning one. $2,400$, $20,000$, and $200,000$ represent the number of images used for in-context learning and fine-tuning, respectively.
In-context
Fine-tuning
$\mu AP$ $2,400$
$2,400$
$20,000$
$200,000$
Novel patterns $56.65$ $41.15$ $60.01$ $67.52$
Base patterns $83.56$ $73.44$ $73.25$ $20.06$
In-context
Fine-tuning
$R@1$ $2,400$
$2,400$
$20,000$
$200,000$
Novel patterns $60.86$ $46.70$ $65.73$ $77.44$
Base patterns $85.85$ $78.04$ $77.77$ $27.99$
6.6 Ablation Studies
Table 5: The comparison between different visual prompting methods for incorporating image-replica pairs. Our ImageStacker not only achieves the highest performance but also maintains efficient training and inference. ‘Infer’ and ‘Train’ are in ‘$10^{-3}s/img$’ and ‘$s/iter$’, respectively.
  Method
$\mu AP$
$R@1$
Infer
Train
@@ 6.6 Ablation Studies
ImageStacker outperforms common visual prompting methods. Models can integrate image-replica pairs at both feature and image levels. At the feature level, we append an extra self-attention layer to the last block of ViT to enable interaction between the class token of the replica and the patch tokens or class token of the image-replica pairs (see Fig. 8 (2) and (3)). At the image level, besides our proposed ImageStacker, we can directly add the image-replica pairs to the replica (See Fig. 8 (4)) or concatenate the images along the height (width) dimension (see Fig. 8 (5)). We compare these different designs in Table 5, drawing three main observations: (1) Incorporating image-replica pairs at the feature level is ineffective, resulting in about a $4\sim 6\%$ performance drop in $\mu AP$. This is attributable to insufficient interaction between the replica and its image-replica pair. (2) Directly adding image-replica pairs to the replica significantly degrades the performance. This is because imposing a strong priori restriction without specific meaning can be detrimental. (3) Both ImageCombiner and ImageStacker significantly improve performance. Compared to ImageCombiner, ImageStacker achieves a greater performance gain ($+8.18\%$ in $\mu AP$ and $+7.52\%$ in $recall@1$), while requiring only about $1/4$ inference workload and $1/3$ training workload. Also, ImageStacker enhances the baseline by $13.81\%$ in $\mu AP$ and $13.00\%$ in $recall@1$, with nearly the same efficiency.
  Method $\mu AP$ $R@1$ Pattern Acc.
Basline (Any) $42.84$ $47.86$
Random $50.42$ $54.83$ $30.11$
Approximate $56.65$ $60.86$ $62.90$
Accurate $56.68$ $60.78$ $64.60$
Ground truth $56.99$ $60.96$ $100.00$
Zero Shot $49.71$ $53.76$
Lower Bound $28.55$ $32.25$
Upper Bound $99.25$ $99.84$
The proposed pattern retrieval method is effective. Table 6 outlines several methods for obtaining the image-replica pair for a given query: (1) Random - selecting an image-replica pair randomly from the entire image-replica pool. (2) Approximate retrieval - using the ImageStacker model. (3) Accurate retrieval - employing another model. (4) Ground truth - using pattern ground truth of queries (not available in practice). ImageStacker can also be applied in a zero-shot setting by duplicating the query two times as the pseudo-prompt (Zero Shot). The lower bound is obtained by using an incorrect image-replica pair, while the upper bound is achieved by using the query and its original image as the image-replica. Their pattern retrieval accuracy is shown in the next section. The demonstration of these methods is visualized in the Fig. 9. We observe that: (1) our approximate retrieval method achieves performance comparable to that of the accurate method ($-0.03\%$ $\mu AP$), and even comparable to using the pattern ground truth ($-0.34\%$ $\mu AP$). It significantly outperforms the random method ($+6.23\%$ $\mu AP$). (2) The zero-shot setting surpasses the baseline by $6.87\%$ in $\mu AP$ and $5.90\%$ in $recall@1$, demonstrating the generalizability of our method. (3) It is non-trivial that using the query itself and its original image as the image-replica nearly achieves $100\%$ performance (see Table 3: without our AnyPattern, using the query itself and its original image can only improve the $\mu AP$ from $16.18\%$ to $18.11\%$), providing evidence of the emergence of in-context learning again.
Our pattern retrieval method achieves high accuracy. This section shows the accuracy of our pattern retrieval methods achieve. The accuracy is defined as
$acc\ =\ \frac{1}{N}\sum_{i=1}^{N}acc_{i}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\#\left(P_{q_{i}}\cap P_{s_{i}}\right)}{{}\#\left(P_{q_{i}}\right)}\right),$ (6)
@@
The pattern retrieval accuracy and the corresponding performance for the four retrieval methods are displayed in Table 6. Our observations are as follows: (1) compared to random selection, our method of approximate pattern retrieval shows an improvement of $+32.79\%$ in $acc$, leading to a $+6.23\%$ improvement in $\mu AP$; (2) while using ground truths can achieve $100\%$ accuracy in pattern retrieval, there is little to no room for further improvement in $\mu AP$. These findings indicate that, although our pattern retrieval is not perfect, it is relatively sufficient to retrieve an example pair to condition the feed-forward process.
Figure 12: The failure cases of our method include: (a) the example pair is incorrect, as in case (1), where the query contains the Pyramid pattern, while the example contains Mosaic and AutoSeg patterns; and (b) the presence of numerous visual distractors, as in case (4), where some references depict scenarios that are inherently similar to the original image. The original image is highlighted in green, while incorrect matches are indicated in red.
Utilizing multiple image-replica pairs per query further improves the performance. We also discover that leveraging multiple image-replica pairs for a given query can further enhance performance. For each query, we retrieve its top-$N$ image-replica pairs and repeat the $9$-channel ImageStacker $N$ times. The final similarity between a query and a gallery is then calculated as the maximum value among $N$ similarities. The performance change in relation to the number of image-replica pairs is depicted in Fig. 10. Compared to using only one image-replica pair, employing $10$ image-replica pairs boosts the $\mu AP$ to $59.59\%$ ($+2.94\%$) and $recall@1$ to $63.16\%$ ($+2.30\%$).
Our method remains effective even with fewer image-replica pairs in the pool. Currently, the AnyPattern pool contains 10 image-replica pairs for each pattern combination. As shown in Fig. 11, we find that for our ImageStacker, one image-replica pair per pattern combination is sufficient to improve performance. This implies that when we utilize only one image-replica pair per query for inference, 10 image-replica pairs per pattern combination in the pool achieves similar performance to just one pair. This further demonstrates the practicality of our method: when operators of a deployed ICD system identify novel patterns, they need only manually add one example pair to the pool.
6.7 Failure Cases
As shown in Fig. 12, we conclude two failure cases: (1) Our method may fail when it retrieves incorrect example pairs. This is reasonable because in-context learning fundamentally relies on the example pairs to “learn”; and (2) Our method may fail when there are many visual distractors. The hard negative problem is indeed a longstanding issue for ICD. To further enhance performance, future work could focus on addressing these two types of failure cases.
@@ 7 Conclusion
This paper considers a practical scenario, i.e. in-context Image Copy Detection (ICD). Unlike the standard updating of ICD, in-context ICD aims to prompt a trained model to recognize novel-patterned replication using a few example pairs, without requiring re-training. To advance research on in-context ICD, we present AnyPattern, a dataset featuring $100$ tampering patterns. We further propose ImageStacker, a method that directly stacks an example pair onto a query along the channel dimension. The stacking design conditions (modifies) the query feature and thus enables a better matching for the novel patterns. Experimental results highlight the substantial performance improvement gained by AnyPattern and ImageStacker. We hope our work draws research attention to the critical real-world problem in ICD systems, i.e., the fast reaction against novel patterns. Beyond the problem of ICD, we also explore the value of AnyPattern in identifying style mimicry by text-to-image diffusion models.
Limitations and future work. Although training ImageStacker with AnyPattern significantly improves performance on novel patterns, it still falls short compared to base patterns. Based on AnyPattern, future work may focus on developing more effective and efficient in-context learning methods to reduce overfitting on base patterns and close this performance gap.
Appendix
@@ Appendix A Demonstration of the AnyPattern dataset
This section shows the $100$ patterns utilized in the creation of the AnyPattern dataset. The original image is presented in Fig. 13; the generated replicas are in the following tables, with the names of the training patterns indicated in black, and the names of the test patterns depicted in blue. The majority of the constructed patterns includes a degree of randomness while a minority of the patterns remain constant. To demonstrate the variability, each pattern is replicated four times for a single image. To our knowledge, the assembled AnyPattern is the most extensive pattern set currently available. The online version of this demonstration is available at https://huggingface.co/datasets/WenhaoWang/AnyPattern/viewer.
@@
Figure 13: The original image that used to add $100$ different patterns
@@
Randomly shift, swap, or invert the channel of an image (Papadakis and Addicam, 2021). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly enhance the edge contrast of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly rearrange (shuffle) the pixels within an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Replicate the effect of a kaleidoscope on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Simulate the effect of viewing an image through a medium that randomly distorts the light. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly simulate the effect of zooming in or out. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply a random Fourier Domain Adaptation (Yang and Soatto, 2020) to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Dividing two random images into patches and then interweave them. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Random use PCA (Krizhevsky et al., 2012) to alter the intensities of the RGB channels in an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply random non-linear distortions within a grid to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Random perceptual group pixels in an image (Neubert and Protzel, 2014). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly change an image with RGB space to HSV, HLS, LAB, YCrCb, LUV, XYZ, or space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly simulate the effect of ”melting” pixels down the image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly replace the style of an image with another one’s (Zhang and Dana, 2017). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly transfer an image to animation style (Chen et al., 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly perform adversarial attacks to an image (Kim, 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply Contrast Limited Adaptive Histogram Equalization (Reza, 2004) to equalize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply DropOut (Srivastava et al., 2014) or DropBlock (Ghiasi et al., 2018). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply WaveBlock (Wang et al., 2022) on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply Wavelet (Torrence and Compo, 1998) with random frequency to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Randomly change the bright and dark areas of the image by manipulating its tone curve. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Erase a random area of an image (Zhong et al., 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Apply AutoAugment (Cubuk et al., 2018) to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Use the Segment Anything (Kirillov et al., 2023) model to preform segmentation. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Remove the haze from an image using dark channel prior (He et al., 2010). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]